Eunice is 2 years less than 3 times her brother's age. Four years from now, Eunice will be twice as old as her brother. How old are they now?
Answer by mananth(12270)

Brother's age be x
Eunice's age = 3x-2

after 4 years
Brother will be x+4
Eunice will be (3x-2)+4= 3x+2
3x+2 = 2( x+4)
3x+2=2x+8
x=6 brother's age
Eunice age = 3x-2==> 3*6-2=16 years